OpenStudy (anonymous):

Simplify the expression. -5+i / 2i Any help would be amazing, thank you!

OpenStudy (danjs):

(-5 + i) / (2i ) i guess, maybe not, and are those just i's or root of -1

OpenStudy (danjs):

the way it is typed in is \[-5 + \frac{ i }{ 2 } * i\]

OpenStudy (anonymous):

@DanJS I'm sort of new to this site so I don't know how to type out equations in the proper format but, but it's -5 + i over 2i. Like: -5+i _____ 2i I hope that makes sense! And I think it's the root of -1. :)

OpenStudy (solomonzelman):

If like that, I would excpet that i does denote \(\sqrt{-1}\). What I would do is multiply times i on top and bottom to rationalize the denominator.

OpenStudy (danjs):

are the notifications not working on the site at the moment?

OpenStudy (solomonzelman):

My reply notifications are a bit off, but message notificatios are on....

OpenStudy (danjs):

yeah, you don't want imaginary i on the bottom, recall that i^2 = -1 and a complex number is in the form a + bi , a - bi

OpenStudy (danjs):

\[\frac{ -5 + i }{ 2i } = \frac{ -5 }{ 2i } + \frac{ i }{ 2i }\] the i cancels in the second term, and need to multiply the first term by i / i to get rid of that i in the bottom

OpenStudy (anonymous):

I think I'm doing it wrong...can you check? @DanJS So I tried following the steps you gave me and right now I'm at: \[\frac{ -5i }{ 2 } + \frac{ 2 }{ 1 }\] Am I doing it right? :(

OpenStudy (anonymous):

Sorry that I'm kind of clueless. You could say math isn't exactly my strongest subject. :p

OpenStudy (danjs):

sorry, i am not receiving notifications at all

OpenStudy (danjs):

\[\frac{ -5 }{ 2i } * \frac{ i }{ i } + \frac{ 1 }{ 2 }\]

OpenStudy (danjs):

remember i^2 = -1 or i = root(-1) That denominator in the first term is 2*i *i = 2*i^2 = 2* -1 = -2

OpenStudy (danjs):

\[\frac{ 5i }{ 2 } + \frac{ 1 }{ 2 }\]

OpenStudy (danjs):

that is it, if you want it in the normal form of a complex number a+bi, rearrange a bit \[\frac{ 1 }{ 2 }+\frac{ 5 }{ 2 }i\]

OpenStudy (anonymous):

I think I understand it now. You helped me a lot! Thank you so much, @DanJS!